OpenStudy (anonymous):

Please help! I don't get arithmetic sequences! What is the sum of the finite arithmetic series? (-10)+0+10+20+...+130

hartnn (hartnn):

Do you know what a common difference is ? and how to find it ?

OpenStudy (anonymous):

No, I don't. THAT'S what I'm trying to figure out. Do I just keep adding them together or subtract?

hartnn (hartnn):

In arithmetic sequence, when you add a certain number to a term, you get the next term. example : 2,5,8,11,.... here if i add 3 to each term, i would get the next term! does this make sense ?

OpenStudy (anonymous):

Yes, so far.

hartnn (hartnn):

(-10)+0+10+20+...+130 what would i add to each termof this sequence to get the next term ?

OpenStudy (anonymous):

-10 right? Because -10+0=-10.

hartnn (hartnn):

sure ? one of the term is 10, you would add +10 or -10 to get the next term 20 ?

OpenStudy (anonymous):

+10. So I just keep adding +10 until I get the next term (+130)?

hartnn (hartnn):

this was only for you to find the "common difference" d of the arithmetic sequence! so if you add +10 to each term, you get the next term, thats why your common difference d = 10 got this so far ?

OpenStudy (anonymous):

Yes. So far, so good.

hartnn (hartnn):

cool :) now the sum, There is a ready-made sum formula for arithmetic sequence which you can use \(\Large S_n = (n/2) [2a_1 +(n-1)d]\) or \(\Large S_n = (n/2)[a_1 +a_n]\) where, n = number of terms a1 = 1st term = - 10 here an = last term f = common difference = d =10 here

hartnn (hartnn):

either case , now we just need to find 'n' which is how many terms are there in the sequence, any ideas ?

OpenStudy (anonymous):

Sorry I'm taking so long. I'm trying to figure it out on paper too. (a+10) (a+2(10)) is how i wrote it out to solve it right?

OpenStudy (anonymous):

So how exactly do i find 'n'? I'm totally lost!!!

hartnn (hartnn):

there are 2 ways, you can find 'n' mentally... 10 is 3rd term 20 is 4th term 30 is 5th term and so on... continuing this pattern, 100 will be 12th term! 110 will be 13th term 120 will be 14th term and lastly 130 will be 15th term makes sense ?

hartnn (hartnn):

the other way is to use the n'th term formula, \(a_n = a_1 +(n-1)d\) but if you got that above way to find 'n', lets not go into this

OpenStudy (anonymous):

Ooooooooh! I've got it so far!

OpenStudy (anonymous):

and is that other way more difficult?

hartnn (hartnn):

excellent! so now we have all the terms of that sum formula. and no, the other method is also easy! i will tell you other method to get 'n' after we finish finding sum

hartnn (hartnn):

use any one of the sum formulas n = 15 a1 = 1st term = -10 an = last term = 130 d = common difference = 10

hartnn (hartnn):

the 2nd formula is easier, you can try that first

OpenStudy (anonymous):

Okay, this is what I have so far: sn=n(a1+an)/2 s15=15(-10+130)/2

hartnn (hartnn):

going good :)

OpenStudy (anonymous):

Okay I got: s15=900. Did I do that right?

hartnn (hartnn):

very good! 900 is the correct sum! :D

hartnn (hartnn):

now for that other way of finding 'n' we know last term = 130, an = 130 d = 10 a1 = 1st term = -10 formula : \( \large a_n = a_1 + (n-1)d \\ \large 130=-10+(n-1)10 \\ 130+10 = 10 (n-1) \\ 140 = 10 (n-1)\\ 14 = n-1 \\ n = 14+1 =15 \) thats how you get number of terms = 15

OpenStudy (anonymous):

Ooooooooh! I think I like this way better! I'll have to practice it more for future reference! :) Thank you very much!

hartnn (hartnn):

most welcome ^_^ you can go through this link for better understanding of sequences :) http://openstudy.com/study#/updates/503bb2a0e4b007f9003103b0

OpenStudy (anonymous):

Thanks again! Have an awesome night!

hartnn (hartnn):

you too :) and have the sweetest dreams :)